OpenStudy (anonymous):

The polar coordinates of a point are given. Find the rectangular coordinates of each point. a. (5,(π / 4)) b. (-2,(π / 6))

OpenStudy (freckles):

\[\cos(\theta)=\frac{x}{r} \text{ or } r \cos(\theta)=x \\ \text{ and } \\ \sin(\theta)=\frac{y}{r} \text{ or } r \sin(\theta)=y\]

OpenStudy (freckles):